OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x Answers: cos x - 4 cos x sin^2x -sin^3x + 2 sin x cos x -sin^2x + 2 sin x cos x 2 sin^2x cos x - 2 sin x cos x

OpenStudy (mathmate):

try using the identity cos(A+B)=cos(A)cos(B)-sin(A)sin(B) sin(A+B)=sin(A)cos(B)+cos(A)sin(B) You may need to go in two steps: cos(2x+x)=... and substitute cos(2x)=cos(x+x)=... and sin(2x)=sin(x+x)=...